OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through (4, 2) and (-2, -3)?

OpenStudy (anonymous):

5x- 6y- 8 = 0?

OpenStudy (anonymous):

y=5/6 x - 4/3..... arrange it.....

OpenStudy (anonymous):

y=mx+c is the slope-intercept form